A ball is thrown vertically upward from the ground with an initial speed of 80 ft/sec. Its height is given by h = 80t - 16t2. How high does the ball go? When does the ball hit the ground?
Answer by funmath(2873)

A ball is thrown vertically upward from the ground with an initial speed of 80 ft/sec. Its height is given by h = 80t - 16t2. How high does the ball go? When does the ball hit the ground?
When it's on the ground h=0.
0=80t-16t^2 (factor)
0=8t(10-2t) (set each part = to 0)
8t=0
8t/8=0/8
t=0 is when you throw it.
10-2t=0
10-10-2t=0-10
-2t=-10
-2t/-2=-10/-2
t=5 s is when it hits the ground.****
Halfway between those two times is when it's at its peak. t=5/2=2.5
h=80(2.5)-16(2.5)^2
h=200-16(6.25)
h=200-100
h=100 ft Is as high as the ball will go.***
Happy Calculating!!!